Explain the following terms:


• Principal quantum number

• Azimuthal quantum number

• Magnetic quantum number

• Spin quantum number

Grade:12

1 Answers

Gaurav Sharma
15 years ago

 

QUANTUM NUMBERS
Quantum numbers describe the position of specific electrons on the energy level diagram. Quantum numbers are of four types;-


Principal quantum number


This number represents the name, size and energy of the shell to which the electron belongs. Different values of n mean different energy levels.
 The value of ‘n’ lies between 1 to ∞.
Value of n           =          1   2  3  4  5  6
Shell                    =          K  L  M  N  O  P
• As higher the value of ‘n’ greater is the distance of the shell from the nucleus.
• Higher the value of ‘n’ greater is the value of energy.
• The maximum number of e- possible in a level is 2n2


Azimuthal Quantum number


It describes the spatial arrangement of electron cloud and angular momentum. It also gives the name of the subshell associated with the main shell.
It  is denoted by ‘l’and its values lies between 0,1,2,...(n-1). Different values of l mean diff sublevels. In a sublevel all the e- have nearly the same energy. Within any level, the lowest energy sublevel is s, then p, then d, then f.
 l = 0           s - subshell     l = 3  f - subshell
l = 1            p – subshell   l = 4  g - subshell
l = 2           d – subshell

Magnetic Quantum Number


This quantum number is designated by the symbol ‘m’. It describes the orientation of electron cloud. For each value of ‘l’ the magnetic quantum number ‘m’  have integral values from –l to +l including zero i.e., total of 2l+1
For l = 0  m = 0
if l = 1 (p orbital), m= -1, 0, 1
if l = 2 (d orbital) m = -2, -1, 0, 1, 2
one orientation corresponds to one orbital thus for l=1 there is 3 orientations which are designed as px,py,pz.